OpenStudy (anonymous):

y+6=-3(x-1) what is the slope

OpenStudy (zale101):

Your equation is not simplified yet. So, simplify it first until it becomes y=mx+b where m is the number that represents the slope

OpenStudy (anonymous):

Subtract 6 from both sides and use distributive property. For example: Y + 4 = 2(x + 9) First distributive property Y+4= 2x + 18 Than to isolate y subtract 4 from both sides and you get Y= 2x + 14 In a y= mx + b situation M= the slope and b is the y int. if this helps